The boíling point of liquid hydrogen is 20. 3k at at atmospheric pressure. What is it`s temperature on the Celsius and Fahrenheit scales

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

The temperature in^{\circ}Cwould be20.3-273=-252.7^{\circ}C

The temperature inFwill be(-252.7\times \dfrac{9}{5})+32=-422.8F